OpenStudy (anonymous):

what is the LCM of 224 and 430?

OpenStudy (anonymous):

Prime Factorization of 224 = 2*2*2*2*2*7 Prime Factorization of 430 = 2*5*43 Therefore, the LCM is 2*2*2*2*2*5*7*43 = 48160